# Conformal Depression Prediction

### v1

@@ 5       $\hat{y}_{i}=\hat{f}(X_{i})$
7             if $\hat{y}_{i}\in[\frac{m-1}{M}K,\frac{m}{M}K)$ then
@@ 8                   $B_{m}\cup\{y_{i}\}$
       // Calculate subinterval according to Eq 7
22       $[y_{m,lo},y_{m,hi}]=E_{m}(q,\mathcal{P}_{m})$
23       if $\hat{y}_{n+1}\in B_{m}$ then
24             $\mathcal{C}_{\alpha}(X_{n+1})=[y_{m,lo},y_{m,hi}]$
@@
26
27 end for
28
29return $\mathcal{C}_{\alpha}(X_{n+1})$
Algorithm 2 Depression Recognition with CDP-ACC
@@ IV EXPERIMENTS
IV-A Data
In order to explore uncertainty quantification methods in depression recognition, we work on two classic facial depression datasets, AVEC 2013 [47] and AVEC 2014 [48]. AVEC 2013 is a subset of the AViD-Corpus, consisting of 150 facial videos. The videos range in length from 20 to 50 minutes, with a frame rate of 30 FPS and a resolution of 640 * 480 pixels. The content of the videos includes facial recordings of subjects during human-computer interaction tasks, such as reading assigned content and improvising on given sentences. We divide the dataset into 70 videos for training, 30 videos for validation (calibration), and 50 videos for testing. The facial depression levels are annotated using BDI-II scores [7], with each video assigned a single value representing the depression level. The BDI-II questionnaire consists of 21 questions, each with multiple-choice answers scored from 0 to 3. The total BDI-II score ranges from 0 to 63: 0-13 indicates minimal depression, 14-19 indicates mild depression, 20-28 indicates moderate depression, and 29-63 indicates severe depression.
AVEC 2014, a subset of AVEC 2013, consists of two tasks: Northwind and Freeform. It comprises 150 Northwind videos and 150 Freeform videos, totaling 300 videos. The original videos of AVEC 2014 are similar to those used in the AVEC 2013 but include five pairs of previously unseen videos, replacing a few videos considered unsuitable. We use 200 videos for training, 50 videos for validation (calibration set), and 50 videos for testing.
It is ensured that the three datasets had similar distributions in terms of age, gender, and depression severity. There is no overlap between the three datasets. Depression severity labels are still based on BDI-II scores [7]. The calibration set should meet the exchangeability with the test set [53]. The better the calibration set, the better the performance of CDP. The videos in the AVEC 2014 dataset are relatively short, with an overlap of 8 frames between adjacent video clips, whereas there is no overlap in the AVEC 2013 dataset. During data processing, landmarks in all video frames are detected using OpenFace [54], and facial region alignment is performed using the eyes, nose, and mouth.
@@ IV-B Experimental Setup
Predictive Model. We focus on post-processing methods and do not have strict requirements on the model. For the purpose of our research without loss of generality, we adopt the classic C3D [55, 19] network as the predictive model for depression recognition. A slight adjustment to the fully connected (FC) layer could better suit the task [20]. In particular, for vanilla regression, we employ MSE Loss as the learning objective during training and modify the last FC layer from 4096 to 64 to enhance training stability and prevent overfitting. In the context of Quantile Regression (QR [49]), we use 99 quantiles evenly spaced between 0.01 and 0.99 for the pinball loss. Then, the output of the last FC layer transforms from a single output in vanilla regression to 99 outputs in QR. To validate the effectiveness of CDP-ACC in the field of depression recognition, we also employ SlowFast [56] as a backbone.
@@
UQ for Depression Recognition. To quantify the uncertainty of depression recognition models, we perform NLL [46] and QR [49] that achieves UQ by retraining the model. We also implement several post-processing methods related to CP [27, 42] to construct valid confidence intervals for depression recognition. Specifically, CQR [43] is a post-processing method that applies conformal prediction to calibrate the interval of pre-trained QR. Its conformal score is $max\{\hat{f}(X;\alpha/2)-y,y-\hat{f}(X;1-\alpha/2)\}$. CDP represents the method for uncertainty quantification in vanilla regression, as shown in Figure 2. Its conformal score is defined as $|y-\hat{y}|$, which achieves marginal coverage. CDP-ACC is the method that achieves approximate conditional coverage and provides valid and adaptive confidence intervals. We set the miscoverage rate $\alpha$ to 0.1. Given that the depression score range is from 0 to 63, we use truncation to limit $\hat{y}_{lo}$ to not be less than 0, and $\hat{y}_{hi}$ not to be greater than 63. For CDP-ACC, we set bins number $M$ to 10 and upper bound $K$ to 63. In addition, we also conduct UQ experiments on seven publicly available general regression datasets (for more information, please refer to the supplementary materials).
@@
Metrics. In depression recognition, we commonly use mean absolute error (MAE) and root mean squared error (RMSE) as evaluate metrics for predicting errors, aiming to quantify the accuracy of the prediction results. To further measure the inherent uncertainty in predictions, we apply the prediction interval coverage probability (PICP) and mean prediction interval width (MPIW) to assess the uncertainty of confidence intervals. PICP and MPIW are two seemingly contradictory metrics. For example, if the PICP worsens, the MPIW may actually improve. We aim to minimize MPIW while ensuring PICP coverage rate. However, if we cannot ensure that PICP is not less than (1 - $\alpha$) $\times$ 100%, merely shortening the interval width is meaningless. Similarly, constructing an interval with a larger width to ensure coverage is also meaningless. For instance, we all know that at 100% confidence level, the interval for the patient’s depression prediction is [0, 64], which is correct but meaningless. With the condition that PICP satisfies $1-\alpha$, a smaller MPIW is considered more desirable.
$PICP=\frac{1}{|\mathcal{D}_{test}|}\sum_{i}^{|\mathcal{D}_{test}|}\mathbf{1}(y_{i}\in\mathcal{C}(X_{i}))\times 100\%,\ \ X\in|\mathcal{D}_{test}|$ (8)
$MPIW=\frac{1}{|\mathcal{D}_{test}|}\sum_{i}^{|\mathcal{D}_{test}|}|\hat{y}_{i,hi}-\hat{y}_{i,lo}|,\ \mathcal{C}(X_{i})=[\hat{y}_{i,lo},\hat{y}_{i,hi}]$ (9)
TABLE I: Results of prediction error
@@
TABLE II: Results of uncertainty quantification
@@
Furthermore, we utilize size-stratified coverage (SSC [57, 44]) to assess the extent to which different methods achieve conditional coverage for subsets of varying depression severity. We divide all possible test data into $G$ bins based on the severity of depression (shown in Table IV), denoted as $B_{1},B_{2},\ldots,B_{G}$. The formula for SSC is as follow:
$SSC=\underset{g\in\{1,..,G\}}{\min}\frac{1}{|\mathcal{I}_{g}|}\sum_{i\in\mathcal{I}_{g}}{\mathbf{1}\{y_{i}\in\mathcal{C}(X_{i})\}}$ (10)
where $\mathcal{I}_{g}\subset\{1,2,\ldots,n_{\text{test}}\}$ represents the observations falling into bin $g$, $|\mathcal{I}_{g}|$ denotes the size of $\mathcal{I}_{g}$ and $\mathbf{1}(\cdot)$ denotes indicator function.
@@ IV-C Results
In this section, we provide a comprehensive evaluation of uncertainty quantification methods for depression recognition through experimental results and analysis. Firstly, Table I displays the prediction errors for vanilla regression, NLL and QR, allowing for a comparison of their accuracy in depression recognition. Furthermore, we compare the uncertainty quantification performance of different UQ methods in depression recognition (shown in Table II), demonstrating the superiority of CDP-ACC. We present in Table III the approximation level of different methods for conditional coverage. We also demonstrate the uncertainty analysis of CDP-ACC for some samples with considerable errors to further enhance insight into CDP-ACC (shown in Table V). Additionally, in the supplementary material, we plot confidence intervals for different methods, providing an intuitive way to discern the strengths and weaknesses of each method.
TABLE III: Results of conditional coverage on AVEC 2014
@@
Table I shows the prediction errors of the C3D and SlowFast under different datasets and methods. We can find that comparing vanilla regression using MSE or NLL as the learning objective. The prediction error for quantile regression is slightly smaller in AVEC 2013. This indicates the 0.5 quantile value of QR has a smaller error than the vanilla regression and NLL. However, in the experiments with the SlowFast model in AVEC 2014, the NLL yields smaller errors compared to the other two methods. This also indicates that if only prediction errors are considered, the NLL method as well as the QR may present better performance. However, both methods demonstrate instability in training processes and results in depression recognition.
Table II shows the application of several uncertainty quantification methods in different depression recognition models. Due to NLL and QR heavily relying on the training process, it’s intuitive to observe that the coverage of QR is the worst, followed by NLL. In the current context of depression recognition, inevitable overfitting, unstable training process and imbalanced datasets make it difficult to obtain effective quantile values or fitting Gaussian distributions. For pre-trained QR, CQR is a valid method to calibrate coverage by adjusting the interval width. In this way, CQR can provide adaptive confidence intervals, but its adaptiveness primarily stems from QR, while the conformal prediction mainly provides theoretical guarantees for marginal coverage. In situations where QR exhibits overconfidence, CQR unavoidably relies on average marginal coverage. As for CDP-ACC, it is an approximate condition coverage method for vanilla regression. It can reduce the interval width as much as possible by providing adaptive confidence intervals while maintaining coverage.
Table III displays the conditional coverage levels of different UQ methods (with $\alpha=0.1$) for various levels of depression severity, which are categorized into four groups (shown in Table IV). Therefore, a SSC closer to $1-\alpha$ indicates better conditional coverage of the UQ method for different levels of depression severity. We can clearly observe that NLL and QR perform the worst in terms of SSC due to unstable training and overconfidence. As an effective method for quantifying depression uncertainty, CDP can only guarantee marginal coverage, hence its results are not as good as CQR. However, the conditional coverage of CQR is limited by the pre-trained QR model. In contrast, CDP-ACC can directly approximate conditional coverage based on the predictions of vanilla regression models, which does not require retraining the QR model and achieves better conditional coverage.
TABLE IV: The Relation Between BDI-II and Depression Severity Level
@@
TABLE V: depression recognition with uncertainty quantification
$\mathbf{Label\ y}$ 11 11 33 43
$\mathbf{C3D\ \hat{y}}$ 0.56 18.4 19.9 25.1
$\mathbf{Interval}$ [0,11] [11,33] [11,33] [10,43]
In the Table V, we present several specific examples with notable errors in depression predictions. Solely relying on the depression prediction of the model, we cannot acquire the model’s confidence in the prediction (e.g., whether the model correctly recognized depression). CDP-ACC, a post-processing method, provides valid confidence intervals for the model’s predictions. A wider interval indicates potentially lower confidence in the model’s predictions. In addition, it also provides us with insights into assessing potential risks of depression.
In the context of depression recognition, we require theoretically guaranteed, statistically rigorous uncertainty quantification methods. Through experimentation, we have observed that methods such as retraining models using NLL and QR tend to exhibit overconfidence. Therefore, we expect post-processing methods and propose CDP for uncertainty quantification by providing confidence intervals. Additionally, we have noticed that CQR can calibrate pre-trained QR and provide adaptive prediction intervals. However, the majority of depression recognition methods are currently based on vanilla regression. Inspired by adaptive prediction sets in classification, we introduce CDP-ACC. By achieving approximate conditional coverage, CDP-ACC offers a plug-and-play uncertainty quantification method for pretrained regression-based depression predictor.
@@ V CONCLUSION
The lack of statistically rigorous uncertainty quantification methods results in unreliable depression recognition. By introducing conformal prediction, we develop a plug-and-play uncertainty quantification method CDP to produce confidence intervals for depression recognition. The intervals predicted by CDP can achieve marginal coverage guarantee at any miscoverage rate. Furthermore, akin to the adaptive prediction sets in classification tasks, we introduce CDP-ACC for depression recognition. It alleviates the issue of marginal coverage practices focusing solely on the average interval width while overlooking individuals with great uncertainty by achieving approximate conditional coverage, and provides valid and adaptive confidence intervals for better uncertainty quantification in depression recognition. We expect that this work will attract more researchers to pay attention to the issue of uncertainty quantification in depression recognition.
V-A Limitations
CDP still faces some challenges in practice:
1. Exchangeable calibration set: The effectiveness of CDP relies entirely on the assumption that the calibration set and test samples are exchangeable, swapping test samples with samples from a calibration set does not change the distribution of data in the calibration set. In practical usage, while we can use a validation set that is independent and identically distributed (IID) with the test set as the calibration set, achieving IID in many cases is challenging. Once there is a distribution shift in the test set, the effectiveness of CDP may diminish.
2. Interval width remains wider than expected: CDP can only provide valid intervals as a post-processing method and does not impact the training process. Currently, the error in depression recognition tasks remains considerable, resulting in wider intervals, indicating less confidence in the model’s results.
V-B Ethical Statement
As a post-processing method, CDP does not compromise user privacy. It is important to note that uncertainty quantification in depression recognition needs to be conducted with user consent, and there should be a thorough explanation of the potential for model prediction failures and the possibility of underlying depression risks.
Acknowledgments
This work was funded in part by the National Natural Science Foundation of China under Grants 61972046.

### v3

@@ 5       $\hat{y}_{i}=\hat{f}(X_{i})$
7             if $\hat{y}_{i}\in[L+\frac{m-1}{M}(U-L),L+\frac{m}{M}(U-L))$ then
@@ 8                   $B_{m}\cup\{y_{i}\}$
23       $N_{m}=|B_{m}|$
       // Calculate conformal score according to Eq. (9) and Eq. (10)
24       $\{s_{i}(m,\alpha)\}_{i=1}^{N_{m}}=\{|E_{m}(1-\alpha,y_{i})|\}_{i=1}^{N_{m}}$
25
      // Calculate the ($1-\alpha$)-th quantile of the conformal scores
26       $\hat{s}(m,\alpha)=\mathcal{Q}(\{s_{i}(m,\alpha)\}_{i=1}^{N_{m}},1-\alpha)$
       // Calculate the interval according to Eq. (12)
27       $[y_{m,l},y_{m,h}]=\mathcal{A}(m,\alpha)$
28
29 end for
30if $\hat{y}_{test}\in B_{m}$ then
31       $\mathcal{C}_{\alpha}(X_{test})=[y_{m,l},y_{m,h}]$
32       return $\mathcal{C}_{\alpha}(X_{test})$
@@ 31       $\mathcal{C}_{\alpha}(X_{test})=[y_{m,l},y_{m,h}]$
Algorithm 2 Depression Prediction with CDP-ACC
@@ IV EXPERIMENTS
IV-A Dataset
In order to verify the effectiveness of our uncertainty quantification methods in depression prediction, we conduct experiments on two commonly used facial depression datasets, AVEC 2013 [48] and AVEC 2014 [49]. AVEC 2013 is a subset of the AViD-Corpus, consisting of 150 facial videos. The videos range in length from 20 to 50 minutes, with a frame rate of 30 FPS and a resolution of 640 $\times$ 480 pixels. The content of the videos includes facial recordings of subjects during human-computer interaction tasks, such as reading assigned content and improvising on given sentences. We divide the dataset into 70 videos for training, 30 videos for validation (calibration), and 50 videos for testing. The facial depression levels are annotated using BDI-II scores [6], with each video assigned a single value representing the depression level. The BDI-II questionnaire consists of 21 questions, each with multiple-choice answers scored from 0 to 3. The total BDI-II score ranges from 0 to 63: 0-13 indicates minimal depression, 14-19 indicates mild depression, 20-28 indicates moderate depression, and 29-63 indicates severe depression, as shown in Table IV.
AVEC 2014, a subset of AVEC 2013, consists of two tasks: Northwind and Freeform. It comprises 150 Northwind videos and 150 Freeform videos. The original videos of AVEC 2014 are similar to those used in the AVEC 2013 but include five pairs of previously unseen videos, replacing a few videos considered unsuitable. We use 200 videos for training, 50 videos for validation (calibration), and 50 videos for testing.
The videos in the AVEC 2014 dataset are relatively short, with an overlap of 8 frames between adjacent video clips, whereas there is no overlap in the AVEC 2013 dataset. During data processing, landmarks in all video frames are detected using OpenFace [56], and facial region alignment is performed using the eyes, nose, and mouth.
@@ IV-B1 Prediction Model
This work focuses on methods for quantifying uncertainty in depression prediction rather than on the predictive performance of the models themselves. We use the classic C3D [57, 19] and the SlowFast [58] networks as baseline models for facial depression prediction. In particular, for vanilla regression, we employ MSE loss as the learning objective during training and modify the last FC layer from 4096 to 64 to enhance training stability and prevent overfitting. In the context of Quantile Regression (QR [50]), we use 99 quantiles evenly spaced between 0.01 and 0.99 for the pinball loss. The output of the last FC layer transforms from a single output in vanilla regression to 99 outputs in QR.
IV-B2 UQ for Depression Prediction
@@
To quantify the uncertainty of depression prediction models, we perform NLL [47] and QR [50] that achieves UQ by retraining the model. We also implement several post-processing methods related to CP [26, 44] to construct valid confidence intervals for depression prediction. Specifically, CQR [45] is a post-processing method that applies conformal prediction to calibrate the interval of pre-trained QR. Its conformal score is set as $max\{\hat{f}(X;\alpha/2)-y,y-\hat{f}(X;1-\alpha/2)\}$. CDP represents the method for uncertainty quantification in vanilla regression, as shown in Fig. 2. Its conformal score is set as $|y-\hat{y}|$, which achieves marginal coverage. CDP-ACC is a method that achieves approximate conditional coverage and provides valid and adaptive confidence intervals. We set the miscoverage rate $\alpha$ to 0.1. Given that the depression score ranges from 0 to 63, we use truncation operator to ensure the predictive interval $[\hat{y}_{l},\hat{y}_{h}]\subseteq[0,63]$. For CDP-ACC, we set the subintervals number $M$ to 14, lower bound $L$ to 0 and upper bound $U$ to 63. To validate the generalizability of our method, we also evaluate our UQ method on seven publicly available datasets for general regression task (for more information, please refer to the Appendix).
IV-B3 Metrics
@@
In depression prediction, mean absolute error (MAE) and root mean squared error (RMSE) are commonly used as the evaluation metrics, aiming to quantify the accuracy performance. To further measure the inherent uncertainty in predictions, we adopt the prediction interval coverage probability (PICP) and mean prediction interval width (MPIW) [51]. PICP and MPIW are two seemingly contradictory metrics. For example, if the PICP worsens, the MPIW may actually improve. We aim to minimize MPIW while ensuring PICP coverage rate. If we cannot ensure that PICP is not less than ($1-\alpha$) $\times$ 100%, merely shortening the interval width is meaningless. Similarly, constructing an interval with a larger width to ensure coverage is also meaningless.
$PICP=\frac{1}{|\mathcal{D}_{test}|}\sum_{i=1}^{|\mathcal{D}_{test}|}\mathbf{1}(y_{i}\in\mathcal{C}_{\alpha}(X_{i}))$ (13)
$MPIW=\frac{1}{|\mathcal{D}_{test}|}\sum_{i=1}^{|\mathcal{D}_{test}|}|\hat{y}_{i,h}-\hat{y}_{i,l}|$ (14)
where $X_{i}\in\mathcal{D}_{test}$ and $[\hat{y}_{i,l},\hat{y}_{i,h}]=\mathcal{C}_{\alpha}(X_{i})$.
TABLE I: Prediction errors of the regression models with different training losses and backbones
@@
TABLE II: Uncertainty quantification performance of different methods
@@
Furthermore, we utilize size-stratified coverage (SSC [59, 46]) to assess the extent to which different methods achieve conditional coverage for subsets of varying depression severity. We divide all possible test data into $G$ bins based on the severity of depression (shown in Table IV), denoted as $\mathcal{I}_{1},\mathcal{I}_{2},\ldots,\mathcal{I}_{G}$.
$SSC=\underset{g\in\{1,..,G\}}{\min}\frac{1}{|\mathcal{I}_{g}|}\sum_{i\in\mathcal{I}_{g}}{\mathbf{1}\{y_{i}\in\mathcal{C}_{\alpha}(X_{i})\}}$ (15)
where $\mathcal{I}_{g}\subset\{1,2,\ldots,n_{\text{test}}\}$ represents the set of sample indices falling into bin $g$, $|\mathcal{I}_{g}|$ denotes the size of $\mathcal{I}_{g}$ and $\mathbf{1}(\cdot)$ denotes the indicator function.
@@ IV-C Results
In this section, we provide a comprehensive evaluation of uncertainty quantification methods for depression prediction. Firstly, we show the prediction errors for vanilla regression, NLL and QR, allowing for a comparison of their accuracy performance in depression prediction. Then, we compare the uncertainty quantification performance of different UQ methods in depression prediction. We also perform uncertainty analysis of CDP-ACC for some instances to gain insight into CDP-ACC. Additionally, in the Appendix, we experimentally evaluate the proposed method on several general regression datasets.
TABLE III: Conditional coverage performance of different methods on AVEC 2014
@@
IV-C1 Results of the UQ
Table I shows the prediction errors of the C3D and SlowFast models under different datasets and training losses. We can find that comparing vanilla regression using MSE or NLL as the learning objective, the prediction error for quantile regression is slightly smaller in AVEC 2013. This indicates the QR with the 50% quantile has a smaller prediction error than the vanilla regression and NLL. However, in the experiments with the SlowFast model on AVEC 2014, the NLL yields smaller errors compared to the other two methods. This also indicates that if only the prediction accuracy is considered, the NLL method as well as the QR may present better performance.
Table II shows the results of several uncertainty quantification methods with different depression prediction models. Due to NLL and QR heavily relying on the training process, it’s intuitive to observe that the coverage of QR is the worst, followed by NLL. In the current context of depression prediction, unstable training process and imbalanced datasets make it difficult to obtain effective quantile values or fitting Gaussian distributions. For pre-trained QR, CQR is a valid method to calibrate coverage by adjusting the interval width. In this way, CQR can provide adaptive confidence intervals, but its adaptiveness primarily stems from QR, while the conformal prediction mainly provides theoretical guarantees for marginal coverage. In situations where QR exhibits overconfidence, CQR unavoidably relies on average marginal coverage. As for our CDP-ACC, it is an approximate condition coverage method for regression. It can reduce the interval width as much as possible by providing adaptive confidence intervals while maintaining coverage.
TABLE IV: BDI-II scores and depression severity levels
@@ IV-C1 Results of the UQ
Figure 6: The prediction intervals for CDP and CDP-ACC on some sampled video clips with 90% confidence ($\alpha$ = 0.1). C3D is used as the depression predictor. To protect privacy, we show facial landmarks of the first frame to represent the facial video clips.
In Fig. 6, we present several facial video samples with varying levels of depression and their prediction uncertainties, where the predictor uses the C3D regression model. From the figure, we can see that although the predictor exhibits different prediction biases on samples with different levels of depression, CDP and CDP-ACC provide valid confidence intervals for the model’s predictions. A wider interval indicates potentially larger uncertainty in the model’s predictions. It can also be seen that compared to CDP, CDP-ACC provides tighter confidence intervals while maintaining coverage. On the other hand, we can observe that even though the predictor may have small prediction biases on some samples (e.g., the second example from the left in Fig. 6), its prediction uncertainty can still be large. This reminds us that in clinical practice, we cannot focus solely on prediction accuracy. From this perspective, our proposed uncertainty quantification scheme offers new insights and perspectives for assessing potential risks in depression prediction.
IV-C2 Evaluation of the Conditional Coverage
Table III shows the conditional coverage of different UQ methods (with $\alpha=0.1$) on test instances with various levels of depression severity, which are categorized into four groups ($G=4$, shown in Table IV). A SSC closer to $1-\alpha$ indicates better conditional coverage of the UQ method. We can clearly observe that NLL and QR perform the worst in terms of SSC due to unstable training and overconfidence. As an effective method for quantifying depression uncertainty, CDP can only guarantee marginal coverage, hence its results are not as good as CQR. However, the conditional coverage of CQR is limited by the pre-trained QR model. In contrast, CDP-ACC can directly approximate conditional coverage based on the point prediction-based vanilla regression models, which does not require retraining a QR model and achieves better conditional coverage.
Figure 7: PICP of CDP-ACC with varying $M$ on AVEC 2014.
Figure 8: MPIW of CDP-ACC with varying $M$ on AVEC 2014.
IV-C3 Parameter Analysis
The effectiveness of CDP-ACC largely depends on the accuracy of conditional distribution estimation. In our histogram estimation, the parameter $M$ affects this accuracy. A smaller $M$ implies a coarser histogram estimation, while a larger $M$ results in a finer estimation but reduces the number of sample points within each subinterval $m$, which can also hinder the effectiveness of the approximation. In our experiments, half of the calibration set is used as the validation set for tuning the parameter $M$. As shown in Figs. 7 and 8, the optimal value of the parameter ($M=14$) can be obtained by tuning on the validation set.
@@ V CONCLUSION
By introducing conformal prediction, we propose a plug-and-play uncertainty quantification method CDP to produce confidence intervals for facial depression prediction. The intervals predicted by CDP can achieve marginal coverage guarantee at any given miscoverage rate. Furthermore, we develop CDP-ACC for facial depression prediction. By approximate conditional coverage, it alleviates the issue of marginal coverage practices focusing solely on the average coverage while overlooking individuals with great uncertainty, and provides valid and adaptive confidence intervals for better uncertainty quantification in facial depression prediction. Although our proposed method is used for facial depression prediction, it is also applicable to various other data modalities. We expect that this work will attract more researchers to pay attention to the issue of uncertainty quantification for automated depression prediction in modern context.
To further validate the effectiveness of our CDP-ACC for general regression tasks, we conducted supplementary experiments on seven publicly available datasets, as also used in [45, 52]. These seven datasets are: Physicochemical Properties of Protein Tertiary Structure (bio) [60], BlogFeedback (blog) [61], Facebook Comment Volume Dataset [62], variants one (fb1) and two (fb2), from the UCI Machine Learning Repository [63], and Medical Expenditure Panel Survey (meps19 - meps21) [64]. Regarding the preprocessing of the datasets and the experimental setup, we followed the settings in [45, 52], with detailed information available on GitHub.
Datasets
The Physicochemical Properties of Protein Tertiary Structure (bio) dataset [60] falls under the subject heading of biology. It consists of data extracted from CASP 5-9 and contains 45,730 decoys ranging in size from 0 to 21 angstroms. This dataset is commonly used for multivariate regression analysis.
The BlogFeedback (blog) dataset [61] belongs to the field of social science. It comprises 60,021 instances, each consisting of features extracted from blog posts and a target variable representing the number of comments the post received within 24 hours. This task is a multivariate regression task commonly used to predict the number of comments a post will receive.
The Facebook Comment Volume Dataset (fb1, fb2) [62] is similar to the blog dataset in that it aims to predict the number of comments on a post. It comprises 40,949 instances, each with 53 features. The dataset includes 5 variants; however, we only utilized the first two variants (fb1 and fb2) in our experiment.
The Medical Expenditure Panel Survey (meps 19 - 21) [64] is a series of large-scale surveys conducted on households, individuals, healthcare providers, and employers across the United States. MEPS serves as the most comprehensive data source on healthcare expenses, utilization, and the extent of medical insurance coverage. In our experiments, we exclusively utilized data from panels 19 to 21.
Implementation Details
Our experimental setup follows [45, 52], where we train a conditional quantile estimator (implemented as a three-layer neural network) to handle the regression task described above. We employ the pinball loss (Eq. 4) as our loss function and ensure that the output quantiles are sorted to prevent crossing.
We train the neural network using the Adam optimizer with a learning rate of 0.02. The output quantiles serve as part of the post-processing calibration for conformal prediction. Unlike SCP [26, 44, 27] and CDP-ACC, which do not rely on quantiles, we typically treat the 50% quantile as the output for vanilla regression. We evaluate using Prediction Interval Coverage Probability (PICP) and Mean Prediction Interval Width (MPIW) as metrics, with a miscoverage rate $\alpha$ set to 0.1. For specific definitions of these evaluation metrics, please refer to our main paper.
Results
We present the uncertainty quantification results of several methods, including SCP [26, 44, 27], QR [50], CQR [45], CHR [52], and CDP-ACC, on the aforementioned general datasets. QR and CQR are introduced in the main paper. CHR is a histogram estimation method that calibrates the QR method by constructing nested intervals and integrating them with conformal prediction.
From Table V, it is clear that SCP generally achieves marginal coverage but tends to construct unnecessarily conservative intervals. In contrast, in general regression tasks, QR maintains normal coverage rates and effectively reduces the mean interval width (However, in depression prediction task, QR exhibits a trend of overconfidence, with quantile values overly concentrated around predictions, leading to lower coverage rates.)
CQR and CHR serve as calibration methods for QR, aiming to minimize interval width while maintaining adequate coverage rates. However, CQR is constrained by the simplicity of QR in constructing conformal scores. In contrast, CHR integrates histogram estimation with conformal prediction, enabling more refined optimization of prediction intervals.
For QR, CQR, and CHR, retraining the QR model is crucial. However, vanilla regression models with point prediction can only allow for obtaining prediction intervals through post-processing methods like SCP. Nevertheless, SCP has significant drawbacks such as fixed interval width and overly conservative intervals (unnecessarily wide). CDP-ACC, as an adaptive post-processing method, can be applied to both QR and vanilla regression models. It achieves a finer approximation of conditional distributions through binning, thereby further reducing interval width. From Table V, we observe that CDP-ACC achieves competitive uncertainty quantification results, demonstrating its effectiveness on general regression tasks..
TABLE V: Results of Uncertainty Quantification.
Methods Data PICP MPIW
SCP [26, 44, 27] bio 90.10% 14.87
QR [50] bio 90.05% 14.53
CQR [45] bio 90.15% 14.56
CHR [52] bio 90.35% 13.13
CDP-ACC bio 89.55% 12.26
SCP [26, 44, 27] blog 91.20% 16.10
QR [50] blog 92.35% 14.65
CQR [45] blog 91.20% 14.65
CHR [52] blog 89.85% 10.24
CDP-ACC blog 90.02% 9.56
SCP [26, 44, 27] fb1 88.25% 16.15
QR [50] fb1 92.00% 15.68
CQR [45] fb1 90.00% 15.67
CHR [52] fb1 89.60% 12.23
CDP-ACC fb1 89.15% 12.11
SCP [26, 44, 27] fb2 90.05% 18.51
QR [50] fb2 92.55% 15.44
CQR [45] fb2 89.95% 15.43
CHR [52] fb2 90.25% 12.14
CDP-ACC fb2 89.70% 10.27
SCP [26, 44, 27] meps19 90.05% 32.55
QR [50] meps19 89.80% 31.23
CQR [45] meps19 90.95% 31.37
CHR [52] meps19 90.04% 22.73
CDP-ACC meps19 89.90% 17.00
SCP [26, 44, 27] meps20 88.85% 23.74
QR [50] meps20 88.45% 26.47
CQR [45] meps20 89.10% 26.53
CHR [52] meps20 89.35% 16.56
CDP-ACC meps20 90.05% 16.00
SCP [26, 44, 27] meps21 90.55% 26.58
QR [50] meps21 88.80% 29.91
CQR [45] meps21 90.05% 29.99
CHR [52] meps21 90.05% 19.44
CDP-ACC meps21 91.15% 15.00
${}^{*}\alpha=0.1$